Project

General

Profile

Meeting 2014-11-19

Date: 19th November 2014
Time: 15:00 UK time
Attendance: M Uchida, C Pidcott, J Greis, C Heidt, A Dobbs, R Gamet, C MacWaters, D Adey, E Overton and P Kyberd,
Apologies: None

Agenda and Minutes

0. Actions

  • Create MICE note of channel mappings end-to-end through hardware, electronics and software - DA - Not present
  • Incorporate updates to online plots including check on RF window timing - MU - Ongoing
  • Design a perspex box to protect cryostats - CM _ Done??
  • Check with R. Bayes that this does not affect things in main CDB geometry - C. Heidt - Polished up geometry from C Hunt need to get files from Geoff. - New action GB and MU.
    Adam said V Blackmore wanted info on what is the central fibres number as defined in the Tracker as alpha is always defined with respect to central fibre. When Adam looked he found a variation of 1 in the geometry to what the tracker group thought it should be. To ensure that we understand this C Heidt will compare data with MC and look into it.
  • Produce plot of efficiency vs pt (to see if low pt performance needs improving?) - AD - ongoing probably for quite a while.

1. Software: AD, ES, CH

  • Schedule - AD (DA from now on)
    Passing on to D Adey. All focus so far on meetings next week.

1.1 Documentation - AD, ES, C. Heidt

  • Ongoing.

1.2 Geometry, Calibrations, Configurations, CDB - C. Heidt

  • Only current geometry issue is an action
  • Geometry in CDB is good

1.3 Monte Carlo

  • MC Noise - C. Heidt
    • Look at large cpu load
  • ADC smearing needs fixing - C. Heidt

1.4 Pattern Recognition - AD

  • Find a way increasing error for low pt tracks and implement
    AD Implemented something that he thinks should work but not been tested against residuals - ongoing
  • Does low pt performance need improving from efficiency plots (action 4)?
    Some progress and work ongoing. Maybe move from 4 point track to 5 point track. ~~15% of total tracks are 4 point tracks and there will be a loss of efficiency currently unknown. Analysis issue really.

1.5 Online Displays - MU Ongoing

  • Schedule now to be finished before next CM
  • Incorporate updates to online plots including check on RF window timing - MU Rhys Gardener?

1.6 Performance and Analysis

  • Emittance error - discrepancy by C. Hunt results and ES results (ES ~0.6%, CH ~ 0.8 - 1.5%)
  • Bias difference between trackers to be investigate - C. Hunt, ES
  • Spacepoint efficiency study - C. Heidt, AD
  • Low pt performance - AD
  • Efficiency and Purity figures for different beams and geometries

All ongoing

1.7 Issues

FROM LAST WEEK:

1.8 Tracker Software Paper - AD

Ongoing published as MICE note and will be referred to in MPB report

2. DAQ/Calibration/Controls - DA, EO

  • Tracker DAQ / electronics QA
    • Second DAQ / electronics - Being tested.
  • Plan for integration with beam: doc

Work ongoing and D Adey and Ed Overton and A Bross in attendance. All except cryo 3 (which is 1m away from position due to floor) all vaced down and helium flex line purge being planned. Cooling to begin today or tomorrow.

Next 24 hours very important.

Electronics: EO: all 4 veners under each cryo have been hooked up and tested. One tight cabling issues with cryo 2 but this has been fixed. Now hooked up LBDS cables and integration windows into cryo1 2 for later. US Tracker cryos to be hooked up and reading out first.

DAQ: DA: Manual triggering and "jerry rigged" system set up and being tested to enable other detectors not to be needed.

2.1 MiniDAQ - DA

2.2 Channel Mappings

  • Create comparison plots for channel mapping data if needed - CH -
    Done we think review next week but prob no progress for a while to to EO.
  • Check our channel mapping from end to end and make sure correct - DA, AD, ES, CH
  • Dig up the old documentation for the electronics channel mapping - DA
    • AFE/cassette channel mapping: MICE note 447.

Came from concerns that channels were swapped after cosmics. We think that it is not a problem.

2.3 Controls

  • Create controls system - DA
    Tracker test system working

2.4 Hall Probes - MU

We already have a working system but we will improve the SW and make it more user friendly.

3. LED: EO

  • Polish fibre ends - Imperial
    Done By K Ladhams arrives at RAL today or tomorrow
  • Make documentation mapping LEDs to areas of tracker illuminated - EO
  • Provide documentation to Maurizio about whether it would be helpful to fit LED into TOF to help our timing - EO
    Contacted Maurizio about injecting light into the TOF for trigger

4. Tracker lab - AB, GB, CM
Used and clean.

5. Installation

  • ISIS long shut down: 25th Aug 2014 - 8th Feb 2015
    *Installation Manual _ GB MU
  • Schedule
    • Current schedule at Integration Schedule - timetable still subject to change (constraints from AFC, etc)
    • Add resource loading? - CM
  • External waveguides
    • QA
  • Cryostats - CM, AB, DA
    • Do we want to keep the cryostats cold afterwards?

6. Services in the MICE hall: CM

  • Helium lines
    Water for compressors 100% done, 3 phase power done

7. Step IV Analysis Readiness - MU

  • Tracker alignment - MU - ongoing
  • Step4 commissioning in progress.

8. Risks

8.1 Hardware

  • Daresbury electronics: vacuum, tracker interlocks
  • Reforming external waveguides (could lead to breakage)
  • Northside PRY (not us, but knock effect to us)
  • Power supply

8.2 Software

  • Not yet able to test against multiple tracks per trigger with MC
  • Real data unpacking broken DA

8.3 Tracker Online Emergency Procedure - MU SB

  • Provide documentation for ISIS and eventually shifters in what to do in an emergency situation for the Tracker when running
    Documentation to Include:
    - schematics of the system
    - a list of associated checklists at power on, and who works through them.
    - a list of instructions to shifters on how to operate and monitor the system
    - emergency shutdown instructions
    - power-up instructions
    - trouble-shooting : what to watch for
    - a list of regular maintenance and safety checks

A lot of this is almost certainly going to be a work in progress, but it would be good to get it started. At the very least the emergency shutdown and power on instructions and checklist should be available by end of the year. We'll need a lot of this for the Jan 21st mock data run anyway.

  • First thing to do determine list of could fatally fails. - At face to face meeting in two weeks.

9. AoB and summary of actions.

9.1 AoB

9.2 Revised Actions

  • Check with R. Bayes that this does not affect things in main CDB geometry - C. Heidt - Polished up geometry from C Hunt need to get files from Geoff. - New action GB and MU
  • Ensure central fibre is correctly defined in geometry and with Tracker group and correct if necessary - C Heidt
  • Issues found with NIM paper description of external waveguide end to end mapping check with Geoff and SW - MU
  • installation manual plan - GB MU

Phone Details

  • The four digit company id: 4673 (callers from outside the UK)
  • The pin number: 419357

Phone numbers:

UK (Imperial College --preferred) 0845 545 4460 English
Austria 0820 4000 1503 English
0820 4000 15 02 German
Belgium 070 35 99 45 English
070 35 99 94 Dutch
070 35 98 66 French
Canada 1 712 432 2832 English
1 712 432 2844 French
Germany 01803 001 178 English
01570 248 8303 English (mobile)
01803 001 177 German
01570 248 8101 German (mobile)
Ireland 0818 270 007 English
Italy 848 390 166 English
848 391 819 Italian
Netherlands 0870 001 901 English
0870 001 909 Dutch
Poland 0801 003 543 English
0801 003 533 Polish
South Africa 087 550 0375 English
Spain 902 885 318 English
902 881 200 Spanish
Sweden 0939 2066 400 English
0939 2066 300 Swedish
Switzerland 0848 560 190 English
0848 560 378 French
0848 560 347 German
0848 560 385 Italian
United Kingdom (PowWowNow) 0844 4 73 73 73 English
United States 1 415 363 0833 English

FibreScansTrackMeetUpdate_19_11_14.pdf (35.5 MB) George, Melissa, 19 November 2014 16:31